# BINAQUAL: A Full-Reference Objective Localization Similarity Metric for Binaural Audio

### v1

@@ Abstract
Spatial audio enhances immersion in applications such as virtual reality, augmented reality, gaming, and cinema by creating a three-dimensional auditory experience. Ensuring the spatial fidelity of binaural audio is crucial, given that processes such as compression, encoding, or transmission can alter localization cues. While subjective listening tests like MUSHRA remain the gold standard for evaluating spatial localization quality, they are costly and time-consuming. This paper introduces BINAQUAL, a full-reference objective metric designed to assess localization similarity in binaural audio recordings. BINAQUAL adapts the AMBIQUAL metric, originally developed for localization quality assessment in ambisonics audio format to the binaural domain. We evaluate BINAQUAL across five key research questions, examining its sensitivity to variations in sound source locations, angle interpolations, surround speaker layouts, audio degradations, and content diversity. Results demonstrate that BINAQUAL effectively differentiates between subtle spatial variations and correlates strongly with subjective listening tests, making it a reliable metric for binaural localization quality assessment. The proposed metric provides a robust benchmark for ensuring spatial accuracy in binaural audio processing, paving the way for improved objective evaluations in immersive audio applications.
@@ 1 Introduction
Spatial audio is an advanced auditory technology that creates a three-dimensional sound field, enabling listeners to perceive sounds as originating from specific locations in space. This technology enhances immersion in applications like virtual reality (VR), augmented reality (AR), gaming, and cinematic soundscapes. Binaural audio, a form of spatial audio, utilizes two channels to simulate the physical and perceptual cues involved in sound localization. With smartphones, headphones, and wearable devices emerging as dominant platforms for music, movies, and interactive media, there is a growing interest in immersive auditory experiences that can be delivered conveniently and effectively [1]. In this context, binaural audio is expected to see significant growth, driven by evolving media consumption patterns.
@@
Localization quality assessment refers to the process of quantifying the accuracy and perceptual fidelity of spatial audio reproduction, typically by comparing perceived sound source locations with their intended positions. Traditionally, subjective listening tests such as MUSHRA [3] have been the gold standard for assessing spatial localization quality. However, such tests are time-consuming, expensive, and prone to variability due to individual listener differences and environmental factors [4]. As the demand for reliable and reproducible evaluations grows, effective objective metrics are needed to assess spatial localization quality in diverse contexts.
Developing effective localization quality metrics is a challenging task. Factors such as variabilities in individual listeners, recording environments, and stimulus characteristics can significantly influence the spatial cues in binaural recordings [5]. In recent years, a variety of spatial localization quality metrics have been proposed. Such metrics can be broadly categorized into signal processing-based techniques [6, 7, 8, 9], and data-driven models [10, 11, 12, 13]. However, the robustness of the metrics to variations in binaural recordings is not well understood, primarily due to the lack of rigorous benchmarks. Also, the reproducibility of these metrics is limited, as their source codes are often unpublished.
@@
Third, we investigate whether the metric can distinguish between simulated binaural representations of different surround speaker layouts. Many spatial audio recordings are mixed specifically for distinct surround layouts, such as 5.1 or 7.1. When down-mixed for binaural listening, preserving spatial cues is essential. A metric capable of distinguishing between these layouts ensures that the spatial localization of the audio is accurately preserved during the binaural rendering of multichannel surround-encoded audio.
@@
  • RQ3: Can the metric distinguish between the simulated binaural representations of different surround speaker layouts?
@@ 2.1 Binaural Audio
The realism of binaural audio relies on the precise simulation of spatial auditory cues that humans use to localize sound sources. These cues include interaural time differences (ITD) and interaural level differences (ILD) [16]. ITD refers to the small delay in the arrival time of a sound at one ear compared to the other. ILD describes the difference in sound intensity between the ears. This difference arises due to the head shadowing effect, where the head partially blocks sound waves, reducing their intensity at the farther ear. ITD is most effective for low-frequency sounds, while ILD is more prominent for high-frequency sounds [17].
@@ 2.2 Sound Source Localization
Sound source localization (SSL) is the process of identifying the spatial location of a sound source, including its azimuth, elevation, and distance, based on auditory cues extracted from the sound waves received by a hearing system or sensor array [20]. This process leverages differences in time, intensity, and frequency content between signals captured at multiple points, such as ears or microphones, to determine the source’s location in a three-dimensional space.
The accuracy of SSL in humans is influenced by multiple factors. The first factor is individual listener characteristics such as age, head and ear anatomy, and prior auditory experience. It has been shown that age negatively affects the accuracy of sound source localization in older adults in terms of both azimuth and elevation [21]. Individual differences in head size, ear shape, and torso affect HRTFs, which describe how sounds are filtered by the body before reaching the eardrum. These variations influence localization accuracy [22, 23]. In addition to these, prior auditory experience, such as familiarity with a sound or training in spatial hearing, enhances the ability to interpret localization cues [24].
@@
Lastly, the characteristics of the stimulus also significantly influence the localization accuracy. Makous et al. [28] investigated the influence of sound location on localization error in humans. They found that sound localization is very accurate for sources located directly in front of the listeners where the average error was 2.0° and 3.5° for horizontal and vertical dimensions, respectively. However, as the sound sources were shifted towards peripheral locations, the errors increased significantly, reaching up to 20° at the extreme positions. The impact of the spectral content of the stimulus on the accuracy of sound source localization has also been investigated in several studies. It has been shown that both frequency content and bandwidth of the signal affect the accuracy of localization [29, 30, 31]. High-frequency spectral cues improve elevation judgments and help resolve front-back confusion [32, 33]. Localization accuracy for low-frequency sounds is typically good in azimuth but poorer in elevation due to the lack of spectral cues [32, 34]. Regarding the sound bandwidth impact, previous work showed that wider stimulus bandwidth leads to higher localization accuracy and vice versa [30, 31, 35]. This is because signals encompassing a wider range of frequencies provide richer spatial cues, leading to improved localization accuracy.
@@ 2.4 AMBIQUAL
AMBIQUAL [14] is a full-reference objective metric designed to assess the quality of compressed ambisonic spatial audio. It evaluates both listening quality and localization accuracy directly from B-format ambisonic. More specifically, listening quality is determined from the omnidirectional channel, while localization accuracy is calculated based on a weighted product of similarities from the directional channels in the B-format audio. AMBIQUAL was developed by adapting and extending ViSQOL [38], a full-reference objective metric designed to evaluate the perceptual quality of audio signals. ViSQOL operates by comparing a reference (high-quality) audio signal with a test (degraded) version. For this purpose, it uses the Neurogram Similarity Index Measure (NSIM) [39], which measures the similarity between corresponding time-frequency components of the reference and test signals, providing a quantitative indication of how well the test signal preserves the spectral characteristics of the reference. Unlike ViSQOL, which uses magnitude spectrograms to measure similarities between reference and test signals, AMBIQUAL leverages spectrograms of phase angles (phaseograms). AMBIQUAL quality predictions were compared with the results of a MUSHRA subjective listening test for a varied set of compressed ambisonics contents, and it was shown that AMBIQUAL predictions correlate well with scores from the listening test [14]. However, the AMBIQUAL model is limited to ambisonic content, as it operates on B-format rather than the binaural render. In other words, the model provides a single quality score for an ambisonic recording, regardless of the render angle.
@@ 3 BINAQUAL Model
As mentioned earlier, in AMBIQUAL, localization accuracy is calculated using a weighted product of similarities from directional channels of B-format ambisonics. Given that binaural audio signals consist of only two channels (left and right), BINAQUAL calculates localization similarity as the product of similarities between the reference and test signals in these two channels.
@@
As shown in part (A) of Fig. 1, spectrograms of phase angles (phaseogram) are first created for each channel of reference and test signals. Phaseograms are computed using a 2048-point Short-Time Fourier Transform (STFT) with a 1536-point Hamming window and 50% overlap. To align with gammatone frequency bands derived from the equivalent rectangular bandwidth (ERB) scale, the number of frequency bins is reduced to the first 640. The reference phaseogram is segmented into patches 480 ms long (30 frames, each 16 ms), and each patch is aligned with the corresponding test patch. The similarity between the reference and test phaseogram patches is calculated on a frame-by-frame basis, considering all 640 frequency bins. The resulting NSIM scores are then averaged across 32 frequency bands corresponding to critical gammatone frequency bands. Finally, similarity scores are computed individually for each patch and averaged across all patches to obtain the overall NSIM similarity score for the test signal.
This process is repeated for both channels of the reference and test signals. As shown in part (B) of Fig. 1, the localization similarity score is calculated as the product of the NSIM scores from the two channels. The source code of the BINAQUAL is available online [^1] https://github.com/QxLabIreland/Binaqual.
@@ 4 Evaluation Datasets
To present a comprehensive evaluation of the BINAUQAL model, three evaluation datasets were created. These datasets were generated synthetically by applying specific HRTFs from subject D2 (dummy head) of the SADIE II database [19] to various audio contents using the Binamix Python library [40]. In cases where the desired angle was present in the SADIE database, the actual angle was used to render the content at the desired angle. Otherwise, an interpolation between the nearest angles was performed. The datasets are specifically designed to address research questions one, two, and three, as outlined in Sec. 1. To address research question four, we use the same dataset used in [14]. It is worth mentioning that we do not create a separate dataset for research question five; Instead, the robustness of the model to various contents and the number of point sources is assessed within the scope of research questions one to four. The compiled dataset is called SynBAD (Synthetic Binaural Audio Dataset) and has been made publicly available [^2] https://zenodo.org/records/15431990. A detailed description of each dataset is provided below.
@@ 4.1 Localization Sensitivity Dataset
The localization sensitivity dataset is designed to evaluate the model’s sensitivity to the locations of point sources (RQ1). To generate the dataset, a subset of HRTFs with varying azimuths and elevations from the SADIE II database was selected. The selection was based on the average performance of humans in sound source localization. As mentioned in Sec. 2.2, human localization errors tend to increase as sound sources shift toward peripheral locations. To account for this, the average unsigned localization errors reported in [28] across different azimuth and elevation ranges were used to determine a subset of HRTFs to synthesize binaural samples. Tables A1 and A2 summarize the increments applied for different azimuth and elevation ranges, respectively.
@@
Using the angles and audio contents described above, 14,500 binaural samples with a duration of 2 s and sampling frequency of 48 kHz were synthesized.
@@ 4.3 Surround Layouts Dataset
The surround layouts dataset is designed to evaluate the model’s ability to distinguish between simulated binaural representations of various surround speaker layouts, including 5.1, 5.1.4, 7.1, and 7.1.4 (RQ3). The previous datasets included sources rendered at a large variety of angles. However, the surround speaker dataset contains renders of various contents at only two sets of angles, one with elevation and one without. This is because we will only use this dataset to investigate the sensitivity of the metric to the presence of elevation (i.e., height speaker) in surround speaker setups.
To generate samples of this dataset, for each of the surround layouts mentioned above, two point sources were rendered at the following angles (azimuth, elevation): (60°, 0°) and (300°, 0°), as well as (60°, 50°) and (300°, 30°). We also rendered the sources at the above angles without any speaker layouts to be used as reference signals.
The Binamix library [40] was used to handle the rendering of sources for specific surround speaker layouts. This library uses impulse response interpolation to simulate the VBAP method [42] and render the source at the virtual speaker positions between the discrete angles available for a specific speaker layout. The audio contents described in Table A3 were mixed in different ways to ensure a diverse set of sound source combinations in this dataset. These combinations, along with the purpose of each, have been detailed in Table A4.

### v2

@@ Abstract
Spatial audio enhances immersion in applications such as virtual reality, augmented reality, gaming, and cinema by creating a three-dimensional auditory experience. Ensuring the spatial fidelity of binaural audio is crucial, given that processes such as compression, encoding, or transmission can alter localization cues. While subjective tests are used for evaluating spatial localization quality, they are costly and time-consuming. This paper introduces BINAQUAL, a full-reference objective metric designed to assess localization similarity in binaural audio recordings. BINAQUAL adapts the AMBIQUAL metric, originally developed for localization quality assessment in ambisonics audio format to the binaural domain. We evaluate BINAQUAL across five key research questions, examining its sensitivity to variations in sound source locations, angle interpolations, surround loudspeaker layouts, audio degradations, and content diversity. Results demonstrate that BINAQUAL effectively differentiates between subtle spatial variations and correlates strongly with subjective listening tests, making it a reliable metric for binaural localization quality assessment. The proposed metric provides a robust benchmark for ensuring spatial accuracy in binaural audio processing, paving the way for improved objective evaluations in immersive audio applications.
@@ 1 Introduction
Spatial audio is an advanced auditory technology that creates a three-dimensional sound field, enabling listeners to perceive sounds as originating from specific locations in space. This technology enhances user immersion across applications such as virtual reality (VR), augmented reality (AR), gaming, and cinematic soundscapes. Binaural audio, a form of spatial audio, utilizes two channels to simulate the physical and perceptual cues involved in sound localization. With smartphones, headphones, and wearable devices emerging as dominant platforms for music, movies, and interactive media, there is a growing interest in immersive auditory experiences that can be delivered in a convenient and effective manner [1]. In this context, binaural audio is expected to see significant growth, driven by evolving media consumption patterns.
@@
Localization quality assessment refers to the process of quantifying the accuracy and perceptual fidelity of spatial audio reproduction, typically by comparing perceived sound source locations with their intended positions. While MUSHRA-style protocols [3] are occasionally adapted for evaluating spatial attributes, localization accuracy is more commonly assessed using task-specific methods that directly measure directional perception, such as directional pointing tasks (e.g., Makous et al. [4]) or Minimum Audible Angle thresholds [5]. However, such tests are time-consuming, expensive, and prone to variability due to individual listener differences and environmental factors [6]. As the demand for reliable and reproducible evaluations grows, effective objective metrics are needed to assess spatial localization quality in diverse contexts.
Developing effective localization quality metrics is a challenging task. Factors such as variabilities in individual listeners, recording environments, and stimulus characteristics can significantly influence the spatial cues in binaural recordings [7]. In recent years, a variety of spatial localization quality metrics have been proposed. Such metrics can be broadly categorized into signal processing-based techniques [8, 9, 10, 11], and data-driven models [12, 13, 14, 15]. However, the robustness of the metrics to variations in binaural recordings remains poorly understood, primarily due to the lack of rigorous benchmarks. Also, the reproducibility of these metrics is limited, as their source codes are often unpublished.
@@
Third, we investigate whether the metric can distinguish between simulated binaural representations of different surround loudspeaker layouts. Many spatial audio recordings are mixed specifically for distinct surround layouts, such as 5.1 or 7.1. When down-mixed for binaural listening, preserving spatial cues is essential. A metric capable of distinguishing between these layouts ensures that the spatial localization of the audio is accurately preserved during the binaural rendering of multichannel surround-encoded audio.
@@
  • RQ3: Can the metric distinguish between the simulated binaural representations of different surround loudspeaker layouts?
@@ 2.1 Binaural Audio
The perceived authenticity of binaural audio relies on the accurate simulation of spatial auditory cues that humans use to localize sound sources. Among these, interaural time differences (ITD) and interaural level differences (ILD) play key roles in lateral localization. ITD refers to the small delay in the arrival time of a sound at one ear compared to the other, and is most effective for low-frequency sounds. ILD describes the difference in sound intensity between the ears, primarily due to the head shadowing effect, and is more prominent for high-frequency sounds [18, 19]. However, these are not the only cues involved in spatial perception. For a plausible and perceptually authentic binaural experience, additional factors such as spectral shaping by the outer ear and dynamic cues related to head movement also contribute significantly to sound source localization in elevation and distance [20, 21]
@@ 2.2 Sound Source Localization
Sound source localization is the process of identifying the spatial location of a sound source, including its azimuth, elevation, and distance, based on auditory cues extracted from the sound waves received by a hearing system or sensor array [21]. This process leverages differences in time, intensity, and frequency content between signals captured at multiple points, such as ears or microphones, to determine the source’s location in a three-dimensional space.
Multiple factors influence the accuracy of sound source localization in humans. The first factor is individual listener characteristics such as age, head and ear anatomy, and prior auditory experience. It has been shown that age negatively impacts the accuracy of sound source localization in older adults in terms of both azimuth and elevation [24]. Individual differences in head size, ear shape, and torso affect HRTFs, which describe how sounds are filtered by the body before reaching the eardrum. These variations influence localization accuracy [25, 26]. In addition to these, prior auditory experience, such as familiarity with a sound or training in spatial hearing, enhances the ability to interpret localization cues [27].
@@
Lastly, the characteristics of the stimulus also significantly influence the localization accuracy. Makous et al. [4] investigated the influence of sound location on localization error in humans. They found that sound localization is very accurate for sources located directly in front of the listeners, where the average error was 2.0° and 3.5° for horizontal and vertical dimensions, respectively. However, as the sound sources were shifted towards peripheral locations, the errors increased significantly, reaching up to 20° at the extreme positions. The impact of the spectral content of the stimulus on the accuracy of sound source localization has also been investigated in several studies. It has been shown that both frequency content and bandwidth of the signal affect the accuracy of localization [31, 32, 33]. High-frequency spectral cues improve elevation judgments and help resolve front-back confusion [34, 35]. Localization accuracy for low-frequency sounds is typically good in azimuth but poorer in elevation due to the lack of spectral cues [34, 36]. Regarding the sound bandwidth impact, previous work showed that wider stimulus bandwidth leads to higher localization accuracy and vice versa [32, 33, 37]. This is because signals encompassing a wider range of frequencies provide richer spatial cues, leading to improved localization accuracy.
A wide range of sound source localization models have been proposed [38, 39, 40, 41]. Some of these models, such as those by Lindemann and Baumgartner, are available in the MATLAB Auditory Modeling Toolbox [42, 43].
@@ 2.4 AMBIQUAL
AMBIQUAL [16] is a full-reference objective metric designed to assess the quality of compressed ambisonic spatial audio. It evaluates both listening quality and localization accuracy directly from B-format ambisonic. More specifically, listening quality is determined from the omnidirectional channel, while localization accuracy is calculated based on a weighted product of similarities from the directional channels in the B-format audio.
AMBIQUAL was developed by adapting and extending ViSQOL [46], a full-reference objective metric designed to evaluate the perceptual quality of audio signals. ViSQOL operates by comparing a reference (high-quality) audio signal with a test (degraded) version. For this purpose, it uses the Neurogram Similarity Index Measure (NSIM) [47], which measures the similarity between corresponding time-frequency components of the reference and test signals, providing a quantitative indication of how well the test signal preserves the spectral characteristics of the reference.
Unlike ViSQOL, which uses magnitude spectrograms to measure similarities between reference and test signals, AMBIQUAL leverages spectrograms of phase angles (phaseograms). Phaseograms are segmented into fixed-duration patches, and NSIM is applied locally to each aligned reference and test patch. The final similarity score is obtained by averaging NSIM values across all patches and frequency bands. These scores are aggregated across directional channels to estimate localization fidelity.
AMBIQUAL quality predictions were compared with the results of a MUSHRA subjective listening test for a varied set of compressed ambisonics contents, and it was shown that AMBIQUAL predictions correlate well with scores from the listening test [16]. However, the AMBIQUAL model is limited to ambisonic content, as it operates on B-format rather than the binaural render. In other words, the model provides a single quality score for an ambisonic recording, regardless of the render angle.
@@ 3 BINAQUAL MODEL
As mentioned earlier, in AMBIQUAL, localization accuracy is calculated using a weighted product of similarities from directional channels of B-format ambisonics. Given that binaural audio signals consist of only two channels (left and right), BINAQUAL calculates localization similarity as the product of similarities between the reference and test signals in these two channels. We should note that BINAQUAL is a relative localization similarity measure, not an absolute localization error estimator (i.e., angular degrees).
@@
As shown in part (A) of Fig. 1, spectrograms of phase angles (phaseograms) are created for each channel of the reference and test signals. These phaseograms are computed using a 2048-point Short-Time Fourier Transform (STFT) with a 1536-point Hamming window and 50% overlap. To align the analysis with the perceptually-motivated frequency range of our 32-band gammatone filterbank (50 Hz to 16 kHz), a specific subset of STFT frequency bins is selected. We use the 640 bins corresponding to indices 2 through 641 of the original STFT. At a 48 kHz sampling rate, this selection captures a frequency range from approximately 47 Hz to 15.02 kHz. This range effectively covers the bands of interest from the gammatone filterbank while excluding the DC component (bin 0).
NSIM is then used to compute localized similarity between patch pairs using two components: an intensity term that reflects the similarity of local means, and a structure term that captures the correlation (covariance) between patches. Both components are stabilized with constants derived from the dynamic range, which in the case of phaseograms corresponds to $2\pi$. These frame-wise similarity scores are computed over all 640 frequency bins and then averaged within 32 perceptually motivated gammatone frequency bands. The band-averaged scores are further aggregated (averaged) across time patches to produce a single global similarity score per channel. The final NSIM score ranges from 0 to 1, where 1 indicates identical patches and values closer to 0 indicate low similarity. This bounded nature makes NSIM particularly well-suited for perceptual similarity tasks, offering normalized and interpretable scores, unlike unbounded metrics such as RMSE. Readers are encouraged to refer to [47] for further details about NSIM.
This process is repeated for both channels of the reference and test signals. As shown in part (B) of Fig. 1, the localization similarity score is calculated as the product of the final NSIM scores from the two channels.
The motivation for applying NSIM to phaseograms of binaural audio signals lies in the relationship between phase information and the spatial cues essential for sound localization. As mentioned earlier, ITDs dominate low-frequency localization, while ILDs dominate at high frequencies. These cues are encoded in the temporal and spectral structure of binaural signals. By comparing the phaseograms of left and right channels between reference and test signals, BINAQUAL indirectly assesses the fidelity of the spatial cues, particularly ITD and ILD, embedded in those signals. Since the NSIM metric is sensitive to both local phase distortions and structural mismatches across frequency bands, it provides a perceptually relevant measure of how well the test signal preserves spatial localization characteristics. Thus, high similarity between binaural phaseograms implies preserved localization cues, whereas reduced similarity suggests perceptual degradation in spatial fidelity.
The source code of the BINAQUAL model is available online [^1] https://github.com/QxLabIreland/Binaqual.
@@ 4 EVALUATION DATASETS
To present a comprehensive evaluation of the BINAUQAL model, three evaluation datasets were created. These datasets were generated synthetically by applying specific HRTFs from subject D2 (dummy head) of the SADIE II database [23] to various audio contents using the Binamix Python library [48]. When the desired angle was available in the SADIE database, the corresponding binaural impulse response (BIR) was used directly. For angles not present in the database, suitable BIRs were generated via interpolation at render time. This was achieved by identifying the nearest available angles using a modified Delaunay triangulation and computing interpolation weights based on Euclidean distances. For further details on the interpolation method, readers are referred to [48].
The datasets are specifically designed to address research questions one, two, and three, as outlined in Sec. 1. To address research question four, we use the same dataset used in [16]. It is worth mentioning that we do not create a separate dataset for research question five; instead, the robustness of the model to various contents and the number of point sources is assessed within the scope of research questions one to four. The compiled dataset is called SynBAD (Synthetic Binaural Audio Dataset) and is available online [^2] https://zenodo.org/records/15431990. A detailed description of each dataset is provided below.
@@ 4.1 Localization Sensitivity Dataset
The localization sensitivity dataset is designed to evaluate the model’s sensitivity to the locations of point sources (RQ1). To generate the dataset, a subset of HRTFs with varying azimuths and elevations from the SADIE II database was selected. The selection was based on the average human performance in sound source localization. As mentioned in Sec. 2.2, human localization errors tend to increase as sound sources shift toward peripheral locations. To account for this, the average unsigned localization errors reported in [4] across different azimuth and elevation ranges were used to determine a subset of HRTFs to synthesize binaural samples. Tables A1 and A2 summarize the increments applied for different azimuth and elevation ranges, respectively.
@@
Using the angles and audio content described above, 14,500 binaural samples were synthesized at a sampling rate of 48 kHz. Since the dataset was intended solely for evaluation and not for subjective listening tests, each sample was limited to a duration of 2 s, sufficient for assessing model sensitivity to point source locations.
@@ 4.3 Surround Layouts Dataset
The surround layouts dataset is designed to evaluate the model’s ability to distinguish between simulated binaural representations of various surround loudspeaker layouts, including 5.1, 5.1.4, 7.1, and 7.1.4 (RQ3). The previous datasets included sources rendered at a large variety of angles. However, the surround layouts dataset contains renders of various contents at only two sets of angles, one with elevation and one without. This is because we will only use this dataset to investigate the sensitivity of the metric to the presence of elevation (i.e., height loudspeaker) in surround loudspeaker setups.
To generate samples of this dataset, for each of the surround layouts mentioned above, two point sources were rendered at the following angles (azimuth, elevation): (60°, 0°) and (300°, 0°), as well as (60°, 50°) and (300°, 30°). We also rendered the sources at the above angles without any loudspeaker layouts to be used as reference signals.
The Binamix library [48] was used to handle the rendering of sources for specific surround loudspeaker layouts. This library uses impulse response interpolation to simulate the VBAP method [50], rendering the source at the virtual loudspeaker positions between the discrete angles available for a specific loudspeaker layout. The audio contents described in Table A3 were mixed in different ways to ensure a diverse set of sound source combinations in this dataset. These combinations, along with the purpose of each, have been detailed in Table A4.